# Greedy Heuristics for Sampling-based Motion Planning in High-Dimensional State Spaces

### v1

@@ Abstract
Sampling-based motion planning algorithms are very effective at finding solutions in high-dimensional continuous state spaces as they do not require prior approximations of the problem domain compared to traditional discrete graph-based searches. The anytime version of the Rapidly-exploring Random Trees (RRT) algorithm, denoted as RRT*, often finds high-quality solutions by incrementally approximating and searching the problem domain through random sampling. However, due to its low sampling efficiency and slow convergence rate, research has proposed many variants of RRT*, incorporating different heuristics and sampling strategies to overcome the constraints in complex planning problems. Yet, these approaches address specific convergence aspects of RRT* limitations, leaving a need for a sampling-based algorithm that can quickly find better solutions in complex high-dimensional state spaces with a faster convergence rate for practical motion planning applications.
This article unifies and leverages the greedy search and heuristic techniques used in various RRT* variants to develop a greedy version of the anytime Rapidly-exploring Random Trees algorithm, denoted as Greedy RRT* (G-RRT*). It improves the initial solution-finding time of RRT* by maintaining two trees rooted at both the start and goal ends, advancing toward each other using greedy connection heuristics. It also accelerates the convergence rate of RRT* by introducing a greedy version of direct informed sampling procedure, which guides the sampling towards the promising region of the problem domain based on heuristics. We validate our approach on simulated planning problems, manipulation problems on Barrett WAM Arms, and on a self-reconfigurable robot, Panthera. Results show that G-RRT* produces asymptotically optimal solution paths and outperforms state-of-the-art RRT* variants, especially in high-dimensional planning problems.
keywords:
Sampling-based path planning, Optimal path planning, Informed sampling, Bidirectional search, Greedy heuristics, High-dimensional planning, Reconfigurable robotic
Figure 1: Comparison between relative sizes of the informed sets in planning problems with tortuous solution paths. The $L^{2}$ informed set is estimated based on the current solution cost, therefore, having the larger ellipsoidal area. The $L^{2}$ greedy informed set, on the other hand, only uses the heuristic cost of the state in the current solution with maximum admissible heuristics, thereby reducing the size of the hyperellipsoid.
@@ 1 Introduction
Robot path planning determines a collision-free geometric path from an initial state to a goal state while considering some specific optimizing objectives such as energy or distance travelled (LaValle, 2006). It is a fundamental research area in robotics and many efforts have been made to develop a wide variety of path planning techniques including graph-based searches, artificial potential fields and sampling-based methods (Elbanhawi and Simic, 2014). Unfortunately, it is challenging to find collision-free and optimal paths, especially in continuous search spaces or in high dimensions, and this task is generally known to be a PSPACE-hard computational problem (Reif, 1979).
@@
In order to achieve computational efficiency in high-dimensional continuous state spaces, sampling-based planning algorithms have been proposed, such as probabilistic roadmaps (PRM; Kavraki et al., 1996) and Rapidly-Exploring Random Trees (RRT; LaValle and Kuffner Jr, 2001). Compared to traditional complete graph-search approaches such as Dijkstra’s algorithm (Dijkstra et al., 1959) or A* (Hart et al., 1968), these sampling-based algorithms relax the completeness requirements and find feasible solutions faster in continuous state spaces with probabilistic completeness guarantees, i.e., the probability that the algorithm fails to find the solution, if one exists, decays to zero as the number of samples approaches infinity. Additionally, extensions have been developed to propose anytime favor sampling-based approaches such as PRM* and RRT* (Karaman and Frazzoli, 2011), which not only hold probabilistic completeness guarantees and also have asymptotic optimality property, i.e., the algorithm almost-surely converges towards the optimum as the number of samples approaches infinity.
@@
These anytime sampling-based algorithms are very effective to find optimal solutions on high-dimensional continuous path planning problems. Like RRTs, these approaches rely on random sampling to explore the search over the problem domain (i.e., incrementally expanding the tree through a free space). Unlike RRTs, they extend it by locally optimizing every added vertex (i.e., incrementally rewire the tree during the tree expansion process). This rewiring process makes sure RRT* to asymptotically converge to the optimal solution if enough time is given. However, it is inefficient to rely on random sampling approach especially in large problem domains or high-dimensional state spaces since a lot of computational effort has been wasted on the states that cannot contribute to improve the current solution or exploring into unnecessary regions of the planning domain.
@@
Therefore, several techniques have been investigated to develop extended versions of RRT and RRT*-based algorithms. Ferguson and Stentz (2006) propose an anytime version of RRT by reusing information from previous trees to improve the quality of the current tree and solution path within the allocated time. They realized that limiting the nodes with a lower cost bound to the previous solution cost can increase the probability of finding even better solutions. Such similar sample rejection ideas have been integrated with RRT* to further improve the convergence rate towards the optimum (Akgun and Stilman, 2011; Otte and Correll, 2013). Building upon these concepts, Gammell et al. (2014, 2018) introduced the direct informed sampling technique in their Informed RRT* algorithm, which exploits the current solution path to focus the search only on promising regions of the planning domain for improving the convergence rate of RRT*.
@@
Informed RRT* limits the exploration space of RRT* by directly sampling within a hyperellipsoid determined by the current path cost. This approach indeed outperforms RRT* in terms of convergence rate and final solution quality while maintaining a uniform sample distribution over the informed subset for optimal guarantees. However, the initial solution paths found by the planner might be tortuous, containing a large number of redundant states. Consequently, this results in unnecessarily larger informed sets, slowing down the exploration and convergence process of the algorithm, especially in high dimensions.
@@
This article addresses the computational burden associated with the informed hyperellipsoid proposed in Informed RRT* by introducing a new direct informed sampling procedure. Our approach biases the sampling based on the heuristic information of the states in the current solution path, regardless of the cost, thus mitigating the impact of tortuous initial solution paths. Specifically, we propose a greedy informed set, a subset of the informed set, which greedily exploits the information from the current solution path to rapidly reduce the size of the informed exploration hyperellipsoid, enhancing sampling efficiency (Figure 1).
@@
Applying direct informed sampling techniques to RRT* does indeed improve the convergence rate of the algorithm, as the search is only focused within hyperellipsoids. However, in certain planning scenarios where the initial path takes longer to find, the advantages of focused search become impractical. One way to find faster initial solutions is to use a bidirectional approach. Bidirectional versions of RRT such as RRT-Connect (Kuffner and LaValle, 2000) grows two trees, one rooted in the start and one in the goal, to explore the problem state space and uses a greedy connect heuristic to guide the trees towards each other. Therefore, this article improves the efficiency of initial solution time in planners that use focused sampling techniques by incorporating bidirectional search concepts.
@@
Expanding upon these greedy heuristics, this paper presents Greedy RRT* (G-RRT*), an almost-surely asymptotically optimal sampling-based planning algorithm based on bidirectional search. G-RRT* is presented to demonstrate the benefits of employing greedy search and heuristic techniques in sampling-based planning to achieve practical performance in high-dimensional state spaces. It shows the importance of exploiting search space to improve sample efficiency through the use of a greedy informed set. Additionally, it presents the necessity of maintaining dual rapidly exploring trees in high-dimensional planning to avoid getting stuck in local minima and to quickly find initial solutions (Figure 2).
@@
The benefits of exploiting the problem state space with greedy heuristics are demonstrated through simulated planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, $\mathbb{R}^{8}$ and $\mathbb{R}^{16}$, manipulation problems on Barrett WAM Arms in $\mathbb{R}^{14}$, as well as on a robotic platform, Panthera,—a 5-degree-of-freedom self-reconfigurable mobile robot (Hayat et al., 2019)—in Section 6. All planning problems are tested with the objective of minimizing path length. These experiments show that our approach is capable of finding quick initial solutions with a faster convergence rate and outperforms state-of-the-art methods, particularly in scenarios involving large planning domains.
1.1 Statement of contributions
This article presents the following specific contributions:
  • Proposes a $L^{2}$ greedy informed set which is constructed based on the heuristic information of the current solution path to accelerate the sampling process (Section 3).
  • Presents Greedy RRT* (G-RRT*), a bi-directional asymptotically optimal sampling-based planning algorithm, as an extension of RRT* that leverages on these greedy heuristic concepts (Section 4).
  • Proves the completeness and asymptotic optimality of G-RRT* by building upon findings from the sampling-based planning literature (Section 5).
  • Demonstrates experimentally that the benefits of greedy exploitation in large planning domains yield better success and convergence rates than state-of-the-art algorithms through simulations and real-world robotic platform (Section 6).
@@ 1.1 Statement of contributions
This article is organized as follows. Section 2 describes a path planning problem definition and provides an overview of the sampling-based path planning and background on accelerating RRT* convergence through sampling. Section 3 defines the greedy informed set, which is demonstrated with a G-RRT* algorithm in Section 4. Section 5 analyses the theoretical guarantees of G-RRT*. Section 6 examines and demonstrates G-RRT* in simulations and in real-world experiments with a self-reconfigurable robot, Panthera. Finally, the article concludes with a discussion and future research plans in Section 7.
2 Preliminaries and Background
@@
This section first presents a formal definition of the optimal path planning problem (Section 2.1) and then reviews an existing literature on sampling-based planning algorithms (Section 2.2) and accelerating convergence of these asymptotically optimal planners based on distribution of samples in the state space (Section 2.3).
2.1 Problem Formulation
@@
In robotics, the path planning problem is commonly distinguished into two problems namely: feasible and optimal planning (Karaman and Frazzoli, 2011). The feasible path planning finds a collision-free geometric path from an initial location to a target location. Therefore, there are many possible solutions in feasible planning problems. The optimal path planning, on the other hand, finds a best unique feasible path based on the given optimization objective. The following definitions given in Definition 1 and 2 formally describes the feasible and optimal path planning problems.
Definition 1 ((The feasible path planning problem)).
@@
Definition 2 ((The optimal path planning problem)).
Let $\mathcal{X}\in\mathbb{R}^{n}$ be the state space of the geometric planning problem. Let $\mathcal{X}_{\text{obs}}\subseteq\mathcal{X}$ be the set of all states, $\mathbf{x}$, that are in collision with obstacles, $\mathcal{O}$, and $\mathcal{X}_{\text{free}}=\mathcal{X}\setminus\mathcal{X}_{\text{obs}}$ be the resultant collision-free set. Let $\mathbf{x}_{I}\in\mathcal{X}_{\text{free}}$ be the initial state and $\mathbf{x}_{G}\in\mathcal{X}_{\text{free}}$ be the goal state. Let $\pi:[0,1]\to\mathcal{X}_{\text{free}}$ be a feasible path in continuous domain. The optimal path planning problem is then formally defined as to find a feasible path $\pi^{*}$ in (2) from $\mathbf{x}_{I}$ to $\mathbf{x}_{G}$ through collision-free space that minimizes the given cost objective, $c:\Sigma\to\mathbb{R}_{\geq 0}$,
@@
Formal guarantees provided by sampling-based algorithms concerning these feasible or optimal planning problems are often probabilistic in nature. An algorithm is considered to be probabilistically complete (Definition 3) when, given an infinite number of samples, the probability of finding a solution, should one exist, approaches one. Furthermore, it is described as almost-surely asymptotically optimal (Definition 4), if the same algorithm, with an increasing number of samples, converges asymptotically to the optimum with a probability of one, provided that an optimal solution exists (Karaman and Frazzoli, 2011). Note that almost-sure asymptotic optimality implies probabilistic completeness.
Definition 3 ((Probabilistic completeness)).
@@
Definition 4 ((Almost-sure asymptotic optimality)).
@@
2.2 Sampling-based motion planning
@@
Sampling-based planning algorithms can be classified into multiple-query and single-query. Multiple-query planners construct a graph (roadmap), representing collision-free paths, that can be queried with multiple initial and goal states by computing the shortest path from initial to the goal state. On the other hand, single-query planners only focus on single initial and goal state pairs. The most commonly used algorithms are probabilistic roadmaps (PRMs; Kavraki et al., 1996; Hsu et al., 1998) for multiple-query and rapidly-exploring random trees (RRTs; LaValle and Kuffner Jr, 2001) for single-query applications. These sampling-based planners are often effective in high-dimensional state spaces and save a large amount of computation by avoiding an explicit representation of obstacles in the state space as opposed to deterministic methods such as graph search-based planners. However, these planners only generate feasible solutions (collision-free paths), which may not be guaranteed to be optimal.
@@
In Karaman and Frazzoli (2011), the authors propose new sampling-based multiple and single-query planners, PRM* and RRT*, the variants of PRMs and RRTs, which provide optimality guarantees. RRT* works in a similar way to RRT by incrementally expanding a tree into the free space, guided by randomly sampled states. However, unlike RRT, it takes into account nearby vertices surrounding this randomly sampled state to ensure that the most suitable parent vertex in the tree is selected. Moreover, RRT* also considers the possibility of rewiring nearby vertices with respect to this randomly sampled state. This incremental rewiring process plays a crucial role in guiding RRT* towards asymptotically converging to the optimal solution, a characteristic of paramount importance in sampling-based motion planning research.
@@
The RRT# algorithm (Arslan and Tsiotras, 2013) extends the exploitation process (local rewiring) of RRT* to global rewiring by utilizing dynamic programming methods, resulting in an improvement of the speed of convergence rate. Removal of states that could not improve the current solution will help reduce the time required for convergence. In Karaman et al. (2011), the authors extend the RRT* by introducing the online graph pruning algorithm called branch-and-bound. Here, a vertex in the current tree is deleted if the cost to reach this vertex, plus the lower-bound on the optimal cost to reach the goal (similar to A* admissible heuristic) is more than the cost of the current solution path. These are the vertices that could not improve the current solution, and pruning them does help improve the RRT* convergence time to enable efficient real-time operation. Moreover, modifications have been made to parent selection and rewiring processes of RRT*. Instead of only considering a set of nearby vertices in a hypersphere as in RRT*, Quick-RRT* (Jeong et al., 2019) also considers their ancestry up to a user-defined parameter for possible parent candidates. A similar tree-extending algorithm is introduced in F-RRT* (Liao et al., 2021), which creates random parent vertices near obstacles to find better initial solutions with a faster convergence rate.
@@
Furthermore, various extensions have been proposed to accelerate the convergence rate such as bi-directional versions: RRT-Connect (Kuffner and LaValle, 2000), RRT*-Connect (Klemm et al., 2015), relaxing optimality to near-optimality: SPARS (Dobson and Bekris, 2014), LBT-RRT (Salzman and Halperin, 2016) and much more. Nevertheless, since the sampling process of these approaches employs uniform random distribution over the whole state space, the algorithm’s convergence rate decreases as the number of dimensions increases, requiring longer computation time to generate optimal solutions.
2.3 State-space sampling

### v2

@@ Abstract
Informed sampling techniques improve the convergence rate of sampling-based planners by guiding the sampling toward the most promising regions of the problem domain, where states that can improve the current solution are more likely to be found. However, while this approach significantly reduces the planner’s exploration space, the sampling subset may still be too large if the current solution contains redundant states with many twists and turns. This article addresses this problem by introducing a greedy version of the informed set that shrinks only based on the maximum heuristic cost of the state along the current solution path. Additionally, we present Greedy RRT* (G-RRT*), a bi-directional version of the anytime Rapidly-exploring Random Trees algorithm that uses this greedy informed set to focus sampling on the promising regions of the problem domain based on heuristics. Experimental results on simulated planning problems, manipulation problems on Barrett WAM Arms, and on a self-reconfigurable robot, Panthera, show that G-RRT* produces asymptotically optimal solution paths and outperforms state-of-the-art RRT* variants, especially in high dimensions.
Index Terms:
Sampling-based path planning, Optimal path planning, Informed sampling, Bidirectional search, Greedy heuristics, High-dimensional planning, Reconfigurable robotic.
Figure 1: Comparison between relative sizes of the informed sets in planning problems with tortuous solution paths. The $L^{2}$ informed set is estimated based on the current solution cost, therefore having the larger ellipsoidal area. The $L^{2}$ greedy informed set, on the other hand, only uses the heuristic cost of the state in the current solution with maximum admissible heuristics, thereby reducing the size of the hyperellipsoid.
@@ I Introduction
Robot path planning determines a collision-free geometric path from an initial state to a goal state while considering some specific optimizing objectives such as energy or distance travelled [1]. It is a fundamental research area in robotics, and many efforts have been made to develop a wide variety of path planning techniques, including graph-based searches, artificial potential fields, and sampling-based methods [2]. Unfortunately, it is challenging to find collision-free and optimal paths, especially in continuous search spaces or in high dimensions, and this task is generally known to be a PSPACE-hard computational problem [3].
@@
In order to achieve computational efficiency in high-dimensional continuous state spaces, sampling-based planning algorithms have been proposed, such as probabilistic roadmaps (PRM) [4] and Rapidly-Exploring Random Trees (RRT) [5]. Compared to traditional complete graph-search approaches such as Dijkstra’s algorithm [6] or A* [7], these sampling-based algorithms relax the completeness requirements and find feasible solutions faster in continuous state spaces with probabilistic completeness guarantees, i.e., the probability that the algorithm fails to find the solution, if one exists, decays to zero as the number of samples approaches infinity. Additionally, extensions have been developed to propose anytime favor sampling-based approaches such as PRM* and RRT* [8], which not only hold probabilistic completeness guarantees and also have asymptotic optimality property, i.e., the algorithm almost-surely converges towards the optimum as the number of samples approaches infinity.
@@
These anytime sampling-based algorithms are very effective in finding optimal solutions to high-dimensional continuous path planning problems. Like RRTs, these approaches rely on random sampling to explore the search space of the problem domain (i.e., incrementally expanding the tree through free space). Unlike RRTs, they extend this process by locally optimizing every added vertex (i.e., incrementally rewiring the tree during the expansion process). This rewiring ensures that RRT* asymptotically converges to the optimal solution if given enough time. However, it is inefficient to rely on a random sampling approach, especially in large problem domains or high-dimensional state spaces, as a significant amount of computational effort can be wasted on states that do not contribute to improving the current solution or on exploring unnecessary regions of the planning domain.
@@
Therefore, researchers have investigated several techniques to develop extended versions of RRT and RRT*-based algorithms. Ferguson and Stentz [9] proposed an anytime version of RRT by reusing information from previous trees to improve the quality of the current tree and solution path within the allocated time. They found that limiting nodes with a lower cost bound relative to the previous solution cost can increase the probability of finding better solutions. Such similar sample rejection ideas have been integrated into RRT* to further improve its convergence rate towards the optimum [10, 11]. Building on these concepts, Gammell et al. [12] introduced the direct informed sampling technique in their Informed RRT* algorithm, which exploits the current solution path to focus the search on promising regions of the planning domain, thereby improving the convergence rate of RRT*.
@@
Informed RRT* limits the exploration space of RRT* by directly sampling within a hyperellipsoid determined by the current path cost. This approach significantly outperforms RRT* in terms of convergence rate and final solution quality while maintaining a uniform sample distribution over the informed subset, ensuring optimal guarantees. However, the initial solution paths found by the planner may be tortuous, containing a large number of redundant states. As a result, this leads to unnecessarily larger informed sets, which slow down the exploration and convergence process, especially in high-dimensional spaces.
@@
This article addresses the computational burden associated with the informed hyperellipsoid proposed in Informed RRT* by introducing a new direct informed sampling procedure. Our approach biases sampling based on heuristic information from the states in the current solution path, regardless of their cost, thereby mitigating the impact of tortuous initial solution paths. Specifically, we propose a greedy informed set, a subset of the informed set that greedily exploits information from the current solution path to rapidly reduce the size of the informed exploration hyperellipsoid, enhancing sampling efficiency (Figure 1).
@@
Applying direct informed sampling techniques to RRT* does indeed improve the convergence rate of the algorithm, as the search is focused within hyperellipsoids. However, in certain planning scenarios where the initial path takes longer to find, the advantages of a focused search become impractical. One way to find faster initial solutions is to use a bidirectional approach. Bidirectional versions of RRT, such as RRT-Connect [13], grow two trees—one rooted at the start and one at the goal—to explore the problem state space and use a greedy connection heuristic to guide the trees towards each other. Therefore, this article improves the efficiency of initial solution times in planners that use focused sampling techniques by incorporating bidirectional search concepts.
@@
Expanding on these greedy heuristics, this paper presents Greedy RRT* (G-RRT*), an almost surely asymptotically optimal sampling-based planning algorithm utilizing bidirectional search. G-RRT* is introduced to demonstrate the benefits of employing greedy search and heuristic techniques in sampling-based planning to achieve practical performance in high-dimensional state spaces. It highlights the importance of exploiting the search space to improve sample efficiency through the use of a greedy informed set. Additionally, it shows the necessity of maintaining dual rapidly exploring trees in high-dimensional planning to avoid local minima and to quickly find initial solutions (Figure 2).
@@
The benefits of exploiting the problem state space with greedy heuristics are demonstrated through simulated planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{4}$, $\mathbb{R}^{8}$, and $\mathbb{R}^{16}$; manipulation problems on Barrett WAM arms in $\mathbb{R}^{14}$; and on a robotic platform, Panthera—a 5-degree-of-freedom self-reconfigurable mobile robot [14] (Section VI). All planning problems are tested with the objective of minimizing path length. These experiments show that our approach is capable of finding quick initial solutions with a faster convergence rate and outperforms state-of-the-art methods, particularly in scenarios involving large planning domains.
@@
This article is organized as follows. Section II describes the path planning problem definition and provides an overview of sampling-based path planning, as well as background on accelerating RRT* convergence through sampling. Section III defines the greedy informed set, which is demonstrated with the G-RRT* algorithm in Section IV. Section V analyzes the theoretical guarantees of G-RRT*. Section VI examines and demonstrates G-RRT* in simulations and in real-world experiments with a self-reconfigurable robot, Panthera. Finally, the article concludes with a discussion and future research plans in Section VII.
I-A Statement of Contributions
This article is a continuation of the ideas of the greedy informed set first introduced in our previous work [15], with a focus on formal guarantees, where we analyze the optimality of the greedy informed set, and demonstrates a novel informed sampling-based planning algorithm through extensive experimental evaluations. Specifically, it makes the following contributions:
  • Provides a formal analysis of the greedy informed set, focusing on how exploitation with heuristics affects optimality (Section III, Theorems 1 and 2)
  • Presents Greedy RRT* (G-RRT*), a bi-directional asymptotically optimal sampling-based planning algorithm, as an extension of RRT* that leverages the greedy heuristic concepts (Section IV).
  • Proves the completeness and asymptotic optimality of G-RRT* by building upon findings from the sampling-based planning literature (Section V, Theorem 3).
  • Demonstrates experimentally that the benefits of greedy exploitation in large planning domains yield better success and convergence rates than state-of-the-art algorithms, through simulations and on a real-world robotic platform (Section VI).
II Preliminaries and Background
@@
This section first presents a formal definition of the optimal path planning problem (Section II-A) and then reviews the existing literature on sampling-based planning algorithms (Section II-B) and methods for accelerating the convergence of these asymptotically optimal planners based on the distribution of samples in the state space (Section II-C).
II-A Problem Formulation
@@
In robotics, the path planning problem is commonly distinguished into two problems, namely: feasible and optimal planning [8]. Feasible path planning finds a collision-free path from an initial location to a target location. Therefore, there are many possible solutions in feasible planning problems. Optimal path planning, on the other hand, finds the best unique feasible path based on the given optimization objective. The following definitions given in Definition 1 and 2 formally describe the feasible and optimal path planning problems.
Definition 1 (The feasible path planning problem).
@@
Definition 2 (The optimal path planning problem).
The optimal path planning problem is formally defined as to find a feasible path $\pi^{*}$ in (2) from $\mathbf{x}_{I}$ to $\mathbf{x}_{G}$ through collision-free space that minimizes the given cost objective, $c:\Sigma\to\mathbb{R}_{\geq 0}$,
@@
Formal guarantees provided by sampling-based algorithms concerning these feasible or optimal planning problems are often probabilistic in nature. An algorithm is considered to be probabilistically complete (Definition 3) when, given an infinite number of samples, the probability of finding a solution, should one exist, approaches one. Furthermore, an algorithm is described as almost-surely asymptotically optimal (Definition 4) if, with an increasing number of samples, it converges asymptotically to the optimal solution with a probability of one, provided that an optimal solution exists [8]. Note that almost-sure asymptotic optimality implies probabilistic completeness.
Definition 3 (Probabilistic completeness).
@@
Definition 4 (Almost sure asymptotic optimality).
@@
II-B Sampling-based motion planning
@@
Sampling-based planning algorithms can be classified into multiple-query and single-query categories. Multiple-query planners construct a graph (roadmap) representing collision-free paths that can be queried with multiple initial and goal states by computing the shortest path from the initial state to the goal state. In contrast, single-query planners only focus on individual initial and goal state pairs. The most commonly used algorithms for multiple-query applications are probabilistic roadmaps (PRMs) [4, 16], while rapidly-exploring random trees (RRTs) [5] are typically used for single-query applications. These sampling-based planners are often effective in high-dimensional state spaces and save a significant amount of computation by avoiding an explicit representation of obstacles in the state space, as opposed to deterministic methods such as graph search-based planners. However, these planners only generate feasible solutions (collision-free paths), which may not be guaranteed to be optimal.
@@
Karaman and Frazzoli [8] propose new sampling-based multiple and single-query planners, PRM* and RRT*, the variants of PRMs and RRTs that provide optimality guarantees. RRT* works in a similar way to RRT by incrementally expanding a tree into the free space, guided by randomly sampled states. However, unlike RRT, it takes into account nearby vertices surrounding the randomly sampled state to ensure that the most suitable parent vertex in the tree is selected. Moreover, RRT* also considers the possibility of rewiring nearby vertices in relation to this randomly sampled state. This incremental rewiring process plays a crucial role in guiding RRT* toward asymptotic convergence to the optimal solution, a characteristic of paramount importance in sampling-based motion planning research.
@@
The RRT# algorithm [17] extends the exploitation process (local rewiring) of RRT* to global rewiring by utilizing dynamic programming methods, resulting in an improvement of the speed of convergence rate. Removal of states that could not improve the current solution will help reduce the time required for convergence. Karaman et al. [18] extend the RRT* by introducing the online graph pruning algorithm called branch-and-bound. Here, a vertex in the current tree is deleted if the cost to reach this vertex plus the lower bound on the optimal cost to reach the goal (similar to the A* admissible heuristic) is more than the cost of the current solution path. These are the vertices that could not improve the current solution, and pruning them does help improve the RRT* convergence time to enable efficient real-time operation. Moreover, modifications have been made to parent selection and rewiring processes of RRT*. Instead of only considering a set of nearby vertices in a hypersphere as in RRT*, Quick-RRT* [19] also considers their ancestry up to a user-defined parameter for possible parent candidates. A similar tree-extending algorithm is introduced in F-RRT* [20], which creates random parent vertices near obstacles to find better initial solutions with a faster convergence rate.
@@
Furthermore, various extensions have been proposed to accelerate the convergence rate, such as bi-directional versions like RRT-Connect [13] and RRT*-Connect [21], as well as methods that relax optimality to near-optimality, including SPARS [22] and LBT-RRT [23]. Nevertheless, since the sampling process of these approaches employs a uniform random distribution over the entire state space, the algorithm’s convergence rate decreases as the number of dimensions increases, requiring longer computation times to generate optimal solutions.
II-C State-space sampling